OpenStudy (anonymous):

Simplify 5 cubed root of 162

OpenStudy (anonymous):

is it \[5\sqrt[3]{162}\]

OpenStudy (anonymous):

\[15 \sqrt[3]{6}\]

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

\[5^{3}\sqrt{162} ???\]

OpenStudy (anonymous):

oh yes your first one was right sorry

OpenStudy (anonymous):

\[5\sqrt[3]{162} = 5*\sqrt[3]{2*3*3*3*3} = 5*3*\sqrt[3]{6} = 15\sqrt[3]{6}\]